OpenStudy (ledah):

A rhombus has a short diagonal of 24 cm and a long diagonal of 32 cm. What is the perimeter of the rhombus?

OpenStudy (ledah):

I ended up getting 60 as my best guess.

Parth (parthkohli):



Parth (parthkohli):

You gotta find the hypotenuse of the right triangle that you see right there. Then, you multiply that by 4(because all the sides are equal). Crappy diagram - but you get the point.

OpenStudy (ledah):

If they had a rotate button for the square tool :P

OpenStudy (ledah):

I might post that in feedback... anyways

Parth (parthkohli):

Nope. A rhombus \(\ne\) square, fyi.

OpenStudy (ledah):

Yeah I know but still a rotating button could be useful for diamonds :P

Parth (parthkohli):

Back to the question. You find the hypotenuse by Pythagorean Theorem. \( \color{Black}{\Rightarrow 24^2 + 32^2 = \text{Side of the rhombus}^2}\)

Parth (parthkohli):

A diamond \(\ne\) square, either ;)

OpenStudy (ledah):

You D:< idea ruin-er xD

Parth (parthkohli):

lol :)

OpenStudy (phi):

I would divide both diagonals by 2 before using pythagoras

Parth (parthkohli):

Thanks for reminding that. Whoops.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 12^2 + 16^2 = \text{side of a rhombus}^2}\)

OpenStudy (ledah):

was I supposed to do 12*12?

OpenStudy (ledah):

or 12*2 and 16*2 =80

OpenStudy (phi):

the little 2 in the upper right means multiply by itself twice. so \( 12^2\) = 12*12 so the side of the rhombus squared =12*12+16*16 remember to take the square root to find the side then multiply by 4 to get the perimeter.

OpenStudy (ledah):

ah

OpenStudy (ledah):

okay